OpenStudy (anonymous):

The circumference of a pizza is approximately 44 inches. Estimate the area of the pizza. A. 154 in2 B. 14 in2 C. 7 in2 D. 53 in2

OpenStudy (calculusxy):

The area of a circle equals to: \[r ^{2}\times \pi \]

OpenStudy (calculusxy):

First you need to find the radius of the circle, given the circumference. To do so, we know that 44=2*3.14*r Divide 44/3.14=14.01 Divide 14.01/2=7 the radius is about 7in. use the formula given to find the area of the circle. \[7^{2} \times \pi\] We know that 7 squared=49. Then 49*3.14=153.86 153.86 rounds to 154in. Answer should be A

OpenStudy (anonymous):

thank you! i was working on paper and i just couldent get the right answer